# CT-Mamba: A Hybrid Convolutional State Space Model for Low-Dose CT Denoising

### v4

@@ 3.3.4 Ablation study
To evaluate the effectiveness of the key components of CT-Mamba, we designed three ablation experiments and conducted quantitative validation on patient L506 from the Mayo dataset:
(A1) Effectiveness of deep NPS loss: Excluding deep NPS-loss during training.
(A2) Effectiveness of the Scanning Method: Replacing the Coherence ”Z” Scan with the classic Visual Mamba scanning method.
@@
To further investigate the impact of deep NPS loss of CT-Mamba in terms of noise texture distribution, we conducted an independent NPS analysis of the slice shown in Fig. 8, focusing on the uniform ROI area marked in Fig. 10(a), with a size of 32×32. By comparing the radial NPS characteristics of the ROI area under conditions without deep NPS loss (A1) and with deep NPS-loss (M, complete model), the optimization effect of deep NPS-loss on noise texture distribution can be more intuitively observed. Fig. 10(b) presents the radial NPS comparison of the 32×32 ROI area, showing that the complete model with deep NPS loss aligns more closely with the NDCT image in terms of frequency distribution and amplitude. This indicates that the loss effectively promotes noise suppression and balanced texture distribution, demonstrating the positive impact of our deep NPS loss on noise optimization.
Figure 10: (a) Slice of patient L506 from the Mayo dataset shown in Fig. 8, with the red dashed box indicating the uniform ROI used for NPS analysis. (b) Radial NPS comparison of this ROI, showing the difference in noise distribution between the model without deep NPS loss (A1) and the complete model with deep NPS loss (M).
@@ 4 CONCLUSIONS
In this paper, we propose a hybrid convolutional state-space model, CT-Mamba, for LDCT image denoising. The model combines the multi-scale analysis capability of wavelet transform, the powerful local feature extraction ability of CNN and the long-range dependent modeling advantage of Mamba, so that it can fully capture the details and global information in LDCT images. We propose a new CZSS module, which adopts a spatially coherence “Z” Scan method, which effectively maintains the spatial continuity between adjacent pixels of the image, and further enhances the detail preservation and noise reduction capabilities of the model. Additionally, to the best of our knowledge, this is the first study to use NPS to guide deep learning in LDCT denoising tasks. We designed a deep NPS-loss driven by Mamba, aiming to ensure that the denoised image accurately restores the noise texture distribution of the NDCT image as much as possible, thereby improving the overall image quality and diagnostic value. We evaluated CT-Mamba on two different datasets, and the experimental results show that our model performs excellently in denoising effectiveness, noise texture preservation, and radiomic feature restoration, showing potential to become a representative method of the Mamba framework for LDCT denoising tasks.

### v5

@@ 3.3.4 Ablation study
To evaluate the effectiveness of the core components of CT-Mamba, we designed three ablation experiments and conducted quantitative validation on patient L506 from the Mayo dataset. Each ablation model was trained using data from the remaining nine patients.
(A1) Effectiveness of Deep NPS loss: Excluding Deep NPS loss during training.
(A2) Effectiveness of the Scanning Method: Replacing the Coherence “Z” Scan with the following two methods: (A2.1) The classic row-column scanning strategy adopted by visual Mamba. (A2.2) The serpentine scanning strategy, which traverses alternating directions between adjacent rows or columns.
@@
Figure 11: Comparison of prediction results under three scanning methods: (A2.1) row-column scanning, (A2.2) serpentine scanning, (M) the proposed Z-shaped scanning. The set range of the display window is a window level of 30 HU and a window width of 340 HU (i.e., a range from -140 HU to 200 HU).
Fig. 11 presents the prediction results under three different scanning methods: (A2.1) row-column scanning, (A2.2) serpentine scanning, (M) the proposed Z-shaped scanning. The results demonstrate that the Z-shaped scanning exhibits superior performance in structural modeling. In particular, it significantly enhances the clarity of vascular structures in the liver region, highlighting its advantages in maintaining spatial continuity and improving detail representation.
@@
To further investigate the impact of the Deep NPS loss on noise texture distribution in CT-Mamba, we conducted a radial NPS analysis on a uniform ROI with a size of 32×32, as marked in Fig. 12(a). The corresponding radial NPS of the ROI are shown in Fig. 12(b). By comparing the radial NPS of the ROI under conditions without (A1, orange curve) and with (M, green curve) the Deep NPS loss, the regulatory effect of this loss on noise can be more intuitively observed. The results indicate that the Deep NPS loss effectively guides the model to generate a noise distribution more consistent with that of NDCT. This is reflected in the radial NPS, where both the frequency distribution trend and amplitude response are closer to the ground truth (blue curve), with particularly significant improvements observed in the low-frequency range. This demonstrates that the Deep NPS loss effectively guides the network to produce noise textures more consistent with NDCT, validating its effectiveness in regulating noise structure.
Figure 12: (a) Three representative CT slices from patient L506 in the Mayo dataset, with the red dashed box indicating the uniform ROI used for NPS analysis. (b) Radial NPS comparison of the corresponding ROIs: the green curve represents the complete model with Deep NPS loss (M), the orange curve represents the model without Deep NPS loss (A1), and the blue curve denotes the ground truth.
@@ 4 Conclusions
In this paper, we propose a hybrid convolutional state-space model, CT-Mamba, for LDCT image denoising. The model combines the multi-scale analysis capability of wavelet transform, the powerful local feature extraction ability of CNN and the long-range dependent modeling advantage of Mamba, so that it can fully capture the details and global information in LDCT images. We propose a new CZSS module, which adopts a spatially coherence “Z” Scan method, which effectively maintains the spatial continuity between adjacent pixels of the image, and further enhances the detail preservation and noise reduction capabilities of the model. Additionally, to the best of our knowledge, this is the first study to use NPS to guide deep learning in LDCT denoising tasks. We designed a deep NPS-loss driven by Mamba, aiming to ensure that the denoised image accurately restores the noise texture distribution of the NDCT image as much as possible, thereby improving the overall image quality and diagnostic value. We evaluated CT-Mamba on multiple datasets, and the experimental results show that our model performs excellently in denoising effectiveness, noise texture preservation, and radiomic feature restoration, showing potential to become a representative method of the Mamba framework for LDCT denoising tasks.